OpenStudy (anonymous):

A road running north to south crosses a road going east to west at the point P. Car A is driving north along the first road, and car B is driving east along the second road. At a particular time car A is 10 kilometers to the north of P and traveling at 80 km/hr, while car B is 15 kilometers to the east of P and traveling at 100 km/hr. How fast is the distance between the two cars changing?

OpenStudy (anonymous):

This is just like the last problem.

OpenStudy (anonymous):

Almost exactly.

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

Yes. We can follow the same method.

OpenStudy (anonymous):

I really hate word problems.

OpenStudy (anonymous):

They trip me up

OpenStudy (anonymous):

ok so y=10+100t, x=15+100t

OpenStudy (anonymous):

and dx/dt=100km/hr?

OpenStudy (anonymous):

Almost. Check your speeds.

OpenStudy (anonymous):

oh I didn't see the 80

OpenStudy (anonymous):

y=10+80t

OpenStudy (anonymous):

For reference.

OpenStudy (anonymous):

Yes. Correct. And what is the formula for the distance d?

OpenStudy (anonymous):

d=sqrt(x^2+y^2)

OpenStudy (anonymous):

Mhmm. So what's next?

OpenStudy (anonymous):

take the derivative

OpenStudy (anonymous):

of the distace according x?

OpenStudy (anonymous):

Of the distance according to t

OpenStudy (anonymous):

t?

OpenStudy (anonymous):

Not yet. How about substituting to get an equation for d in terms of t?

OpenStudy (anonymous):

oh yes!

OpenStudy (anonymous):

We have d = sqrt(x^2 + y^2) x = 15 + 100t y = 10 +80t

OpenStudy (anonymous):

yes. simplifying now.

OpenStudy (anonymous):

Good good =) I'm not gonna check your work on this one.

OpenStudy (anonymous):

k so d=sqrt(325+4600t+16400t^2)

OpenStudy (anonymous):

then take dd/dt

OpenStudy (anonymous):

Yup. =)

OpenStudy (anonymous):

You can not substitute also.. \[\frac{dD}{dt} = \frac{d}{dt}[\sqrt{x^2 + y^2}]\] \[=\frac{1}{2\sqrt{x^2 + y^2}}\cdot \frac{d}{dt}(x^2 + y^2)\]\[=\frac{1}{2\sqrt{x^2 + y^2}}\cdot(2x\frac{dx}{dt} + 2y\frac{dy}{dt})\] Makes for a little less work actually.

OpenStudy (anonymous):

I'm confused now.

OpenStudy (anonymous):

Ah. Good point, polpak. Might be simpler for her to understand the other way though.

OpenStudy (anonymous):

Sorry, feel free to ignore me if you like.

OpenStudy (anonymous):

Yeah, what you're doing is right, Nasryn. He just has another method.

OpenStudy (anonymous):

Did you get the derivative? It's still going to have t involved like last time.

OpenStudy (anonymous):

Should get the same answer either way. Your way just required you to integrate then differentiate again. My way is just differentiate then plug/chug.

OpenStudy (anonymous):

yes. dd/dt=(16400t+2300)/sqrt(325+4600t+16400t^2)

OpenStudy (anonymous):

look ridonkulous

OpenStudy (anonymous):

Very good. And what is t?

OpenStudy (anonymous):

0, right?

OpenStudy (anonymous):

I'm not sure.

OpenStudy (anonymous):

I have to go, sorry. I'm going with friends to see some live jazz music =)

OpenStudy (anonymous):

ok no worries.

OpenStudy (anonymous):

But t is 0, so if you plug in 0, that should give you your answer =)

OpenStudy (anonymous):

:) thanks.

OpenStudy (anonymous):

Later.

OpenStudy (anonymous):

\[\frac{dD}{dt} = \frac{1}{2\sqrt{x^2 + y^2}}\cdot(2x\frac{dx}{dt} + 2y\frac{dy}{dt})\] \[\implies \frac{dD}{dt}|_{x=15, y=10} = \frac{1}{2\sqrt{15^2 + 10^2}}\cdot(2(15)(100) + 2(10)(80))\]\[= \frac{1}{2\sqrt{325}}\cdot 2(1500 + 800)\]\[=\frac{2300}{5\sqrt{13}}\]\[={460 \over \sqrt{13}}\]

OpenStudy (anonymous):

:D wonderful thank you.

OpenStudy (anonymous):

You can at least check your answer if you don't follow my method, and/or you can try to figure out how I did it ;) Basically I'm just implicitly differentiating and using the chain rule here.

OpenStudy (anonymous):

yeah, I did it your way because I'm pretty sure that's the way I'm supposed to do it. So thank you!